OpenStudy (anonymous):

The length of the sides of a square measure 2x-5. The length of a rectangle measure 2x, and the width measures x=2.For what value of xis the perimeter of the square the same of the rectangle? Show work F. x=2 H. x=10 G. x=7 I. x=12

OpenStudy (gw2011):

Show perimeter of square = perimeter of rectangle The perimeter of the square=all four sides added together (all the sides are the same) The perimeter of the rectangle=all four sides added together (two of the sides are the same and the other two sides are the same) Therefore, the following is true: 4(2x-5) = 2(2x) + 2(2) 8x-20 = 4x+4 4x = 24 x = 6 Proof: Square: 4(2x-5) = 4(12-5) = 4(7) = 28 Rectangle: 2(12)+4 = 24+4 = 28